OpenStudy (anonymous):

A particle moves along the curve below. y = sqrt(1 + x^3) As it reaches the point (2, 3), the y-coordinate is increasing at a rate of 12 cm/s. How fast is the x-coordinate of the point changing at that instant?

zepdrix (zepdrix):

This is calculus question, yes? Have you learned about derivative shortcuts yet?

OpenStudy (anonymous):

well yeah, however word problems are not my strong suit.

zepdrix (zepdrix):

So our derivative gives us something like this. \[\Large\rm y=\sqrt{1+x^3},\qquad y'=\frac{3x^2}{2\sqrt{1+x^3}}\] Evaluate the derivative at x=2, you're get a numerical value. This represents the slope of the function at the point (2,3). If the instantaneous rate of change of y, \(\Large\rm dy\) is 12 at that point, then what is the instantaneous rate of change of x?\[\Large\rm \frac{dy}{dx}=y'(2)\] \[\Large\rm \frac{12}{dx}=y'(2)\]Hopefully I'm interpreting that correctly...